OpenStudy (anonymous):

solve the equation: 3^z=9sqrt3

OpenStudy (anonymous):

To clarify, is it: \[3^z = 9\sqrt{3}\]

OpenStudy (callisto):

\[3^{z} = 9\sqrt{3}\]\[3^{z} =3^{2}\times3^{(1/2)}\]\[3^{z}=3^{5/2}\]\[z=5/2\]

OpenStudy (anonymous):

yes, tsigma

OpenStudy (callisto):

actually, the question is clear enough and no need to clarify la

OpenStudy (anonymous):

break it up into two parts to make it a little easier. first we are going to deal with 3^z=9 since we know that 3x3 = 9 this part can be simplified to be 3^2 now the sqrt(3) can also be written as 3^1/2 so we put those two together and we get (3^2)(3^(1/2)=9sqrt(3) this simplifies to 3^(5/2)

OpenStudy (anonymous):

Thank you very much--that is very clear and easy for me to understand. :-)

OpenStudy (anonymous):

no prob. that's what we're here for